OpenStudy (anonymous):

is the ordered pair a solution of the given system of equations ? (9,10), 2/3x+5y=56 7x-8/5y=47

OpenStudy (anonymous):

My answer is the ordered is not a solution to the given system

OpenStudy (anonymous):

\[(2/3)(9)+5(10) = 56\]\[6+50=56\]\[56=56\] \[7(9)-((8/5)(10)) = 47\]\[63-16=47\]\[47=47\] so yes they are an ordered pair

OpenStudy (anonymous):

solution

OpenStudy (anonymous):

How did you get 47

OpenStudy (anonymous):

My answer is the ordered is not a solution to the given system

OpenStudy (anonymous):

 Find the soltuion of the systemof equations on each graph,If there is no solution ,state it